# Gravity from (A)dS Yang-Mills: a natural reduction

### v2

@@ Abstract
We investigate the relationship between a one-parameter family of (anti)-de Sitter Yang-Mills models and Einstein-Palatini gravity with matter, realized through Inönü-Wigner contraction of the (A)dS algebra. By setting the gauge group parameter $\alpha$ to zero, the gauge transformation of the potential becomes consistent with the transformation properties of the tetrad form and spin connection. We show that a sector of the Yang-Mills dynamics exists where the gauge fields acquire a geometric interpretation, and the full Yang-Mills symmetry gives rise to diffeomorphisms and local Lorentz transformations. Finally, the corresponding dynamics is consistent with unimodular gravity.
keywords: (anti)-de Sitter gauge theory, first order gravity, unimodular gravity.
@@ 1 Introduction
The search for a formulation of gravity as a gauge theory has a long history. There is an extensive literature on the subject, starting from the seminal papers of Utiyama, Sciama, Kibble, Hayashi, and Nakano [1, 2, 3, 4] and the contributions by Hehl in the 80’s (see for example the reprint [5]), gauging the Lorentz and Poincaré groups. Many of these approaches are reviewed in [6], particularly in the recent textbook [7].
@@
The first step toward a gauge-theoretic formulation of gravity typically consists in moving from the second-order metric formulation of the Einstein–Hilbert action to a first-order, or tetradic-Palatini, formulation [8], which amounts to choosing new, independent degrees of freedom, tetrads and spin connection, and promoting the local Lorentz group to the gauge group. The resulting action does not yield a gauge theory of Yang-Mills type, since the tetrad $e_{\mu}^{a}$ is not a gauge connection, while the spin connection $\omega_{\mu}^{ab}$ is. However, by extending the construction which works in lower dimensions, where the theory is topological [9], the former can be formulated as a gauge theory of $BF$ type [10], provided the so-called simplicity constraint is implemented for the $B$ field [11, 12].
@@
On the other hand, attempts to formulate gravity as a Yang-Mills theory for the whole Poincaré group, so to include the tetrads in the gauge connection, encounter a well known problem: due to the degeneracy of the Cartan-Killing metric of the algebra, the translational sector of the theory is non-dynamical, unless the Yang-Mills action is modified or the gauge group is extended (see for example [13], [14]). In this framework, many existing proposals in the literature exploit the possibility of obtaining the Poincaré group as a Inönü-Wigner contraction of de Sitter or anti-de Sitter groups [15, 16]. A related approach is represented by the MacDowell-Mansouri action  [17] whose geometric interpretation in terms of Cartan geometry is clarified in [18]. For an updated review see [7].
@@
In this paper, we propose yet another formulation of gravity as a one-parameter family of standard Yang-Mills models on Minkowski space for the (anti)-de Sitter group ((A)dS in the following), with the physical content of the theory being recovered in the Inönü-Wigner contraction. The choice of pseudo-orthogonal groups offers two obvious advantages: it allows to encode the would-be geometric fields into components of the extended gauge connection and it guarantees that the scalar product in the algebra, defined via the Cartan-Killing metric, is invariant under the adjoint action of the group and nondegenerate.
@@
One main novelty of our proposal is the invariance in form of the action in the limit $\alpha\rightarrow 0$, being $\alpha$ the contraction parameter from the (A)dS group to the Poincaré group. We shall see that the standard Lagrangian description of the Yang-Mills gauge theory can be applied without modifications, while the geometric interpretation of the gauge fields as tetrad form and Lorentz connection will be recovered as $\alpha\rightarrow 0$, with the Lorentz subgroup as the residual gauge symmetry. The choice of de Sitter or anti-de Sitter group will ultimately be motivated by the fact that they preserve a quadratic form with an appropriately chosen signature, whose four-dimensional reduction is subsequently preserved by the Lorentz group [19].
@@
The idea of using the (A)dS group is not new. Various works have proceeded along similar lines in the past and the current literature. In  [20], the $SO(3,2)$ gauge symmetry is spontaneously broken to the Lorentz group, in such a way as to yield the Einstein-Cartan theory. The Goldstone field associated with the symmetry-breaking mechanism allows the derivation of the gravitational vierbein and spin connection from the original $SO(3,2)$ gauge fields by transitioning to a set of nonlinearly transforming fields through a redefinition that involves the Goldstone field. The work in [21] proposes a non-linear realization of the gauge connection on a fiber bundle, where the fiber is the homogeneous space $SO(1,4)/SO(1,3)$. Then the components of the gauge potential are identified with tetrads and spin connection, leading to Euler-Lagrange equations that have gravitational counterparts in the material sources, linked to curvature and torsion.
@@
More recently, in [22], the Inönü-Wigner contraction, responsible for the transition from the gauge group to the Lorentz group, is related to some mass parameter. Additionally, each gauge field configuration defines an effective geometry through a isomorphism from $\mathbb{R}^{4}$ to a final manifold $\mathbb{M}^{4}$, a deformed space. Another construction is proposed in [23], where a Yang-Mills type gauge action is analyzed, and the gravitational action emerges as a projection of the five-dimensional space on which the de Sitter group fibration is defined. This projection identifies a natural tetrad, which defines the spacetime volume, thereby breaking the original gauge symmetry to $SO(1,3)$. Finally, in [24], gravity models are derived from a topological gauge theory which is a combination of topological invariants of a manifold with an underlying Cartan geometry $G/H$, $G$ being the (A)dS group and $H$ the Lorentz subgroup $SO(1,3)$.
@@
Our analysis starts with a standard Yang-Mills gauge theory on a Minkowski base manifold, associated with a one-parameter family of algebras of pseudo-orthogonal groups, $\mathfrak{g}_{\alpha}$, which recovers the de Sitter, anti-de Sitter, and Poincaré algebras for $\alpha$ being respectively positive, negative, or zero. In the limit where $\alpha$ tends to zero, the structure group reduces to the Poincaré group but the gauge connection is defined in such a way that the action functional does not change, being independent of $\alpha$. However the interpretation of the gauge fields in the action changes radically. In fact, in the limit, the theory is no longer a gauge theory for the whole (A)dS group. The local symmetry of the action reduces to Lorentz making it possible to interpret the emerging model as a gravitational theory. In this work we identify and analyse the set of conditions that must be met for such an interpretation to be well grounded.
@@
The paper is organized as follows. In Section 2 we introduce the Yang-Mills action for the one parameter family of Lie algebras. In Section 2.1, we study the differential identities implied by the Bianchi identity of the Yang-Mills action. This analysis is needed for ensuring consistency with the Bianchi identities of curvature and torsion of gravity in the limit $\alpha\rightarrow 0$. In Section 2.2 we analyse the behavior of the different components of the gauge potential under gauge transformations and we show that, in the limit $\alpha\rightarrow 0$, their transformation properties are consistent with their interpretation as tetrad one-form and Lorentz connection. In Section 3 we derive the Euler Lagrange equations in the presence of matter, and we discuss their formal equivalence with the equations of gravity in the first-order formalism. In Section 4, we complete the identification of tetrads and Lorentz connection in the limit $\alpha\to 0$, by identifying the diffeomorphisms with a sector of the gauge transformations. Finally, we describe the corresponding Euler-Lagrange equations in the limit $\alpha\to 0$, deriving the gravitational equations. We conclude by discussing results and perspectives in Section 5.
@@ 2 The Yang-Mills action for the (A)dS group
with generators $J_{ab},P_{a},a,b=0,\dots 3$ and $K^{cd}_{a,b}=\frac{1}{2}(\delta^{c}_{a}\delta^{d}_{b}-\delta^{d}_{a}\delta^{c}_{b})$. This choice of basis singles out the Lorentz subalgebra spanned by the generators $J_{ab}$ (more details in Appendix A). We introduce a $\mathfrak{g}_{{\alpha}}$-valued connection one-form
$\Omega=\frac{1}{2}\varpi^{ab}J_{ab}+\sqrt{\frac{\lambda}{{{\alpha}}}}\vartheta^{a}P_{a}\,\,\,\in\Omega^{1}(\mathcal{M})\otimes\mathfrak{g}_{{\alpha}}$ (2.2)
with $\varpi_{ab}$ and $\vartheta_{a}$ respectively indicating its components in the Lie algebra generators $J_{ab}$ and $P_{a}$. The local expression on the base manifold $(\mathbb{R}^{(1,3)},\eta)$, with $\eta$ the Minkowski metric, is given by $\Omega=\Omega_{\mu}dx^{\mu}$, $\mu=0,\dots 3$. The parameter $\lambda$ has dimensions of inverse length square and its sign determines the sign of $\alpha$ to ensure that the radicand is positive. [^2] In addition, notice that as a consequence of the dimensional properties of the coefficient $\lambda$, the two components of the connection have different dimensions: $\varpi^{ab}_{\mu}$ has dimensions of the inverse of a length, while $\vartheta^{a}_{\mu}$ is dimensionless.
@@
is naturally decomposed into two components, a Lorentz valued two-form and a second term which is $P_{a}$-valued. We have
@@
The Yang-Mills action for the de (A)dS group reads
$S_{{YM}}=\frac{\hbar}{2g^{2}}\int_{\mathcal{M}}\left(\mathcal{F}\underset{\vbox{\hbox{\scriptsize,}}}{\wedge}*\mathcal{F}\right)\,$ (2.8)
@@
Starting from (2.4), we can explicitly expand the Lagrangian density in (2.8) according to our choice of basis in the algebra. We get a sum of four contributions
$\begin{split}\left(\mathcal{F}\underset{\vbox{\hbox{\scriptsize,}}}{\wedge}*\mathcal{F}\right)=\frac{1}{4}\mathfrak{R}^{ab}\wedge*\mathfrak{R}^{cd}(J_{ab},J_{cd})+\frac{\lambda}{\alpha}\mathfrak{T}^{a}\wedge*\mathfrak{T}^{d}(P_{a},P_{d})+\\ +\frac{\lambda}{2}\mathfrak{R}^{ab}\wedge*\mathfrak{E}^{cd}(J_{ab},J_{cd})+\frac{\lambda^{2}}{4}\mathfrak{E}^{ab}\wedge*\mathfrak{E}^{cd}(J_{ab},J_{cd})\end{split}$ (2.9)
By computing the product in the algebra (see the Appendix in (A.8)), the action finally reads
$S_{YM}=\frac{\hbar}{2g^{2}}\int_{\mathcal{M}}\frac{1}{2}\mathfrak{R}^{ab}\wedge*\mathfrak{R}_{ab}+\lambda\mathfrak{T}^{a}\wedge*\mathfrak{T}_{a}+\frac{\hbar\lambda}{2g^{2}}\int_{\mathcal{M}}\mathfrak{R}_{ab}\wedge*\mathfrak{E}^{ab}+\frac{\lambda}{2}\mathfrak{E}_{ab}\wedge*\mathfrak{E}^{ab}$ (2.10)
The first two terms of the action are similar in form to the Euler classes for the curvature $\mathfrak{R}$ and the two form $\mathfrak{T}$,  [^3] In fact, unlike the Euler classes, which are described using an internal Hodge product, these are characterized by a Hodge product with respect to the metric of the base manifold. while the last two are equivalent in form to the Einstein-Palatini action with cosmological constant [6, 8]. As for the latter contribution, notice that it naturally emerges from the definition of the gauge connection (2.2). Furthermore, by adding a topological term to the Yang-Mills action, defined as
$S_{top}=\frac{\kappa}{16\pi^{2}}\int_{\mathcal{M}}\left(\mathcal{F}\underset{\vbox{\hbox{\scriptsize,}}}{\wedge}\mathcal{F}\right)$ (2.11)
@@
$\begin{split}S_{YM+top}&=\underbrace{\frac{\hbar\lambda}{2g^{2}}\int_{\mathcal{M}}\mathfrak{R}_{ab}\wedge*(\vartheta^{a}\wedge\vartheta^{b})+\frac{\lambda}{2}\vartheta^{a}\wedge\vartheta^{b}\wedge*(\vartheta_{a}\wedge\vartheta_{b})}_{S_{E-P}+{Cosmological\,Constant}}\\ &+\underbrace{\frac{\hbar\lambda}{\gamma g^{2}}\int_{\mathcal{M}}\mathfrak{R}_{ab}\wedge\vartheta^{a}\wedge\vartheta^{b}}_{Palatini\,\,Holst\,\,Term}+\underbrace{\frac{\hbar}{2g^{2}}\int_{\mathcal{M}}\frac{1}{2}\mathfrak{R}_{ab}\wedge*\mathfrak{R}^{ab}+\lambda\mathfrak{T}_{a}\wedge*\mathfrak{T}^{a}}_{{Generalized\,\,Euler-Classes}}\\ &+\underbrace{\frac{\hbar}{4\gamma g^{2}}\int_{\mathcal{M}}\mathfrak{R}_{ab}\wedge\mathfrak{R}^{ab}}_{Pontryagin\,\,Term}+\underbrace{\frac{\hbar\lambda}{2\gamma g^{2}}\int_{\mathcal{M}}\mathfrak{T}_{a}\wedge\mathfrak{T}^{a}-\mathfrak{R}_{ab}\wedge\vartheta^{a}\wedge\vartheta^{b}}_{Nieh-Yan\,\,Term}\end{split}$ (2.12)
Therefore, the addition of a topological term to the Yang-Mills action naturally produces further contributions which can be identified with well known topological invariants of gravity [25, 26, 27, 28]. [^4] For this reason, with an abuse of notation, we name the topological terms in the same way as their counterparts which will be obtained in the contraction limit. The question is whether the formal analogy of the (A)dS Yang-Mills model with first-order Einstein-Palatini gravity can be made physical via a contraction of the structure group to Poincaré, and which conditions should be satisfied for that to happen. To this aim, the choice of basis in the (A)dS algebra and its dependence on $\alpha$ play an essential role. In the limit $\alpha\rightarrow 0$, $\mathfrak{g}_{\alpha}$ reduces to the Poincaré algebra but the action does not change in form and it is still well defined, being independent of the parameter, despite the fact that for $\alpha\to 0$, the connection (2.2) and the curvature (2.4) are ill defined. Therefore the action (2.10) will remain gauge invariant only with respect to the Lorentz group, suggesting a consistent identification with some formulation of gravity if the limit is carefully performed.
@@
    the differential identities implied by the vanishing of the exterior covariant derivative of the curvature two-form, $\mathcal{D}_{\Omega}\mathcal{F}=0$ are consistent with the Bianchi identities of the would be Riemann curvature and torsion;
@@
    the behavior of the gauge fields under the action of the local Lorentz group, in the limit $\alpha\rightarrow 0$, is consistent with their interpretation as tetrad one form and Lorentz connection;

### v3

# Gravity model from (A)dS Yang-Mills theory

@@ Abstract
We investigate the relationship between a one-parameter family of (anti-)de Sitter Yang-Mills models and a model of Einstein-Palatini gravity with matter, realized through Inönü-Wigner contraction of the (A)dS algebra. By setting the group parameter $\alpha$ to zero, the gauge transformation of the potential becomes consistent with the transformation properties of the tetrad form and spin connection. We show that a sector of the Yang-Mills dynamics exists in which the equations decouple. Moreover, a subset of the gauge transformations can be related to diffeomorphisms, leading to the identification of the tetrad field. Finally, the resulting dynamics is consistent with a gravitational dynamics in the first-order formalism.
keywords: (anti-)de Sitter gauge theory, first order gravity
@@ 1 Introduction
The search for a formulation of gravity as a gauge theory has a long history. There is an extensive literature on the subject, starting from the seminal papers of Utiyama, Sciama, Kibble, Hayashi, and Nakano [Utiyama:1956sy, Sciama:1964wt, Kibble:1961ba, 10.1143/PTP.38.491] and the contributions by Hehl in the 1980s (see for example the reprint [Hehl:2023khc]), gauging the Lorentz and Poincaré groups. Many of these approaches are reviewed in [Peldan:1993hi], particularly in the recent textbook [Krasnov:2020lku].
@@
The first step toward a gauge-theoretic formulation of gravity typically consists in moving from the second-order metric formulation of the Einstein-Hilbert action to a first-order, or tetradic-Palatini, formulation [Palatini:1919ffw], which amounts to choosing new, independent degrees of freedom, tetrads and spin connection, and promoting the local Lorentz group to the gauge group. The resulting action does not yield a gauge theory of Yang-Mills type, since the tetrad $e_{\mu}^{a}$ is not a gauge connection, while the spin connection $\omega_{\mu}^{ab}$ is. However, by extending the construction that works in lower dimensions, where the theory is topological [Witten:2007kt], the former can be formulated as a gauge theory of $BF$ type [Birmingham:1991ty], provided the so-called simplicity constraint is implemented for the $B$ field [Plebanski:1977zz, Krasnov:2010olp].
@@
On the other hand, attempts to formulate gravity as a Yang-Mills theory for the whole Poincaré group, so to include the tetrads in the gauge connection, encounter a well- known problem: due to the degeneracy of the Cartan-Killing metric of the algebra, the translational sector of the theory is nondynamical, unless the Yang-Mills action is modified or the gauge group is extended (see, for example, [Kawai], [Aldrovandi:1981mv]). In this framework, many existing proposals in the literature exploit the possibility of obtaining the Poincaré group as a Inönü-Wigner contraction of de Sitter or anti-de Sitter groups [inonu1964group, Inonu]. A related approach is represented by the MacDowell-Mansouri action  [MacDowell:1977jt] whose geometric interpretation in terms of Cartan geometry is clarified in [Wise_2010]. For an updated review see [Krasnov:2020lku].
@@
In this paper, we propose yet another formulation of gravity as a one-parameter family of standard Yang-Mills models on Minkowski space for the (anti) de Sitter group [(A)dS in the following], with the physical content of the theory being recovered in the Inönü-Wigner contraction. The choice of pseudo-orthogonal groups offers two obvious advantages: it allows one to encode the would-be geometric fields into components of the extended gauge connection and it guarantees that the scalar product in the algebra, defined via the Cartan-Killing metric, is invariant under the adjoint action of the group and nondegenerate.
@@
One main novelty of our proposal is the invariance in the form of the action in the limit $\alpha\rightarrow 0$, being $\alpha$ the contraction parameter from the (A)dS group to the Poincaré group. We shall see that the standard Lagrangian description of the Yang-Mills gauge theory can be applied without modifications, while the geometric interpretation of the gauge fields as tetrad form and Lorentz connection will be recovered as $\alpha\rightarrow 0$, with the Lorentz subgroup as the residual gauge symmetry. The choice of de Sitter or anti-de Sitter group will ultimately be motivated by the fact that they preserve a quadratic form with an appropriately chosen signature, whose four-dimensional reduction is subsequently preserved by the Lorentz group [Randono:2010cq].
@@
The idea of using the (A)dS group is not new. Various works have proceeded along similar lines in the past and the current literature. In  [PhysRevD.21.1466], the $SO(3,2)$ gauge symmetry is spontaneously broken to the Lorentz group, in such a way as to yield the Einstein-Cartan theory. The Goldstone field associated with the symmetry-breaking mechanism allows the derivation of the gravitational vierbein and spin connection from the original $SO(3,2)$ gauge fields by transitioning to a set of nonlinearly transforming fields through a redefinition that involves the Goldstone field. The work in [Tseytlin:1981nu] proposes a nonlinear realization of the gauge connection on a fiber bundle, where the fiber is the homogeneous space $SO(1,4)/SO(1,3)$. Then the components of the gauge potential are identified with tetrads and spin connection, leading to Euler-Lagrange equations that have gravitational counterparts in the material sources, linked to curvature and torsion.
@@
More recently, in [Sobreiro:2011hb], the Inönü-Wigner contraction, responsible for the transition from the gauge group to the Lorentz group, is related to some mass parameter. Additionally, each gauge field configuration defines an effective geometry through a isomorphism from $\mathbb{R}^{4}$ to a final manifold $\mathbb{M}^{4}$, a deformed space. Another construction is proposed in [mistretta2023], where a Yang-Mills type gauge action is analyzed, and the gravitational action emerges as a projection of the five-dimensional space on which the de Sitter group fibration is defined. This projection identifies a natural tetrad, which defines the spacetime volume, thereby breaking the original gauge symmetry to $SO(1,3)$. Finally, in [Lazzarini:2024ocr], gravity models are derived from a topological gauge theory which is a combination of topological invariants of a manifold with an underlying Cartan geometry $G/H$, $G$ being the (A)dS group and $H$ the Lorentz subgroup $SO(1,3)$.
@@
Our analysis starts with a standard Yang-Mills gauge theory on a Minkowski base manifold, associated with a one-parameter family of algebras of pseudo-orthogonal groups, $\mathfrak{g}_{\alpha}$, which recovers the de Sitter, anti-de Sitter, and Poincaré algebras for $\alpha$ being, respectively, positive, negative, or zero. In the limit where $\alpha$ tends to zero, the structure group reduces to the Poincaré group but the gauge connection is defined in such a way that the action functional does not change, being independent of $\alpha$. However the interpretation of the gauge fields in the action changes radically. In fact, in the limit, the theory is no longer a gauge theory for the whole (A)dS group. The local symmetry of the action reduces to Lorentz making it possible to interpret the emerging model as a gravitational theory. In this work we identify and analyze the set of conditions that must be met for such an interpretation to be well grounded.
@@
The paper is organized as follows. In Sec. 2 we introduce the Yang-Mills action for the one parameter family of Lie algebras. In Sec. 2.1, we study the differential identities implied by the Bianchi identity of the Yang-Mills action. This analysis is needed for ensuring consistency with the Bianchi identities of curvature and torsion of gravity in the limit $\alpha\rightarrow 0$. In Sec. 2.2 we analyse the behavior of the different components of the gauge potential under gauge transformations and we show that, in the limit $\alpha\rightarrow 0$, their transformation properties are consistent with their interpretation as tetrad one-form and Lorentz connection. In Sec. 3 we derive the Euler Lagrange equations in the presence of matter, and we discuss their formal equivalence with the equations of gravity in the first-order formalism. In Sec. 4, we complete the identification of tetrads and Lorentz connection in the limit $\alpha\to 0$, by identifying the diffeomorphisms with a sector of the gauge transformations. Finally, we describe the corresponding Euler-Lagrange equations in the limit $\alpha\to 0$, deriving the gravitational equations. We conclude by discussing results and perspectives in Sec. 5.
@@ 2 The Yang-Mills action for the (A)dS group
with generators $J_{ab},P_{a},a,b=0,\dots,3$ and $K^{cd}_{a,b}=\frac{1}{2}(\delta^{c}_{a}\delta^{d}_{b}-\delta^{d}_{a}\delta^{c}_{b})$. This choice of basis singles out the Lorentz subalgebra spanned by the generators $J_{ab}$ (more details in Appendix A). We introduce a $\mathfrak{g}_{{\alpha}}$-valued connection one-form
$\Omega=\frac{1}{2}\varpi^{ab}J_{ab}+\sqrt{\frac{\lambda}{{{\alpha}}}}\vartheta^{a}P_{a}\,\,\,\in\Omega^{1}(\mathcal{M})\otimes\mathfrak{g}_{{\alpha}}\$(2.2)
with $\varpi_{ab}$ and $\vartheta_{a}$ respectively indicating its components in the Lie algebra generators $J_{ab}$ and $P_{a}$. The local expression on the base manifold $(\mathbb{R}^{(1,3)},\eta)$, with $\eta$ the Minkowski metric, is given by $\Omega=\Omega_{\mu}dx^{\mu}$, $\mu=0,\dots,3$. The parameter $\lambda$ has dimensions of inverse length square and its sign determines the sign of $\alpha$ to ensure that the radicand is positive [^2] In addition, notice that as a consequence of the dimensional properties of the coefficient $\lambda$, the two components of the connection have different dimensions: $\varpi^{ab}_{\mu}$ has dimensions of the inverse of a length, while $\vartheta^{a}_{\mu}$ is dimensionless..
@@
is naturally decomposed into two components, a Lorentz valued two-form and a second term that is $P_{a}$-valued. We have
@@
The Yang-Mills action for the (A)dS group reads
$S_{{YM}}=\frac{\hbar}{2g^{2}}\int_{\mathcal{M}}\left(\mathcal{F}\underset{\vbox{\hbox{\scriptsize,}}}{\wedge}*\mathcal{F}\right)\,$, (2.8)
@@
Starting from (2.4), we can explicitly expand the Lagrangian density in (2.8) according to our choice of basis in the algebra (see Appendix A). We get a sum of four contributions
$\begin{split}\left(\mathcal{F}\underset{\vbox{\hbox{\scriptsize,}}}{\wedge}*\mathcal{F}\right)=\frac{1}{4}\mathfrak{R}^{ab}\wedge*\mathfrak{R}^{cd}(J_{ab},J_{cd})+\frac{\lambda}{\alpha}\mathfrak{T}^{a}\wedge*\mathfrak{T}^{d}(P_{a},P_{d})+\\ +\frac{\lambda}{2}\mathfrak{R}^{ab}\wedge*\mathfrak{E}^{cd}(J_{ab},J_{cd})+\frac{\lambda^{2}}{4}\mathfrak{E}^{ab}\wedge*\mathfrak{E}^{cd}(J_{ab},J_{cd}).\end{split}$ (2.9)
By computing the product in the algebra [see the Appendix in (A.8)], the action finally reads
$S_{YM}=\frac{\hbar}{2g^{2}}\int_{\mathcal{M}}\frac{1}{2}\mathfrak{R}^{ab}\wedge*\mathfrak{R}_{ab}+\lambda\mathfrak{T}^{a}\wedge*\mathfrak{T}_{a}+\frac{\hbar\lambda}{2g^{2}}\int_{\mathcal{M}}\mathfrak{R}_{ab}\wedge*\mathfrak{E}^{ab}+\frac{\lambda}{2}\mathfrak{E}_{ab}\wedge*\mathfrak{E}^{ab}.$ (2.10)
The first two terms of the action are similar in form to the Euler classes for the curvature $\mathfrak{R}$ and the two form $\mathfrak{T}$,  [^3] In fact, unlike the Euler classes, which are described using an internal Hodge product, these are characterized by a Hodge product with respect to the metric of the base manifold. while the last two are equivalent in form to the Einstein-Palatini action with cosmological constant [Peldan:1993hi, Palatini:1919ffw]. As for the latter contribution, notice that it naturally emerges from the definition of the gauge connection (2.2). Furthermore, by adding a topological term to the Yang-Mills action, defined as
$S_{top}=\frac{\kappa}{16\pi^{2}}\int_{\mathcal{M}}\left(\mathcal{F}\underset{\vbox{\hbox{\scriptsize,}}}{\wedge}\mathcal{F}\right),$ (2.11)
@@
$\begin{split}S_{YM+top}&=\underbrace{\frac{\hbar\lambda}{2g^{2}}\int_{\mathcal{M}}\mathfrak{R}_{ab}\wedge*(\vartheta^{a}\wedge\vartheta^{b})+\frac{\lambda}{2}\vartheta^{a}\wedge\vartheta^{b}\wedge*(\vartheta_{a}\wedge\vartheta_{b})}_{S_{E-P}+{Cosmological\,Constant}}\\ &+\underbrace{\frac{\hbar\lambda}{\gamma g^{2}}\int_{\mathcal{M}}\mathfrak{R}_{ab}\wedge\vartheta^{a}\wedge\vartheta^{b}}_{Palatini\,\,Holst\,\,Term}+\underbrace{\frac{\hbar}{2g^{2}}\int_{\mathcal{M}}\frac{1}{2}\mathfrak{R}_{ab}\wedge*\mathfrak{R}^{ab}+\lambda\mathfrak{T}_{a}\wedge*\mathfrak{T}^{a}}_{{Generalized\,\,Euler-Classes}}\\ &+\underbrace{\frac{\hbar}{4\gamma g^{2}}\int_{\mathcal{M}}\mathfrak{R}_{ab}\wedge\mathfrak{R}^{ab}}_{Pontryagin\,\,Term}+\underbrace{\frac{\hbar\lambda}{2\gamma g^{2}}\int_{\mathcal{M}}\mathfrak{T}_{a}\wedge\mathfrak{T}^{a}-\mathfrak{R}_{ab}\wedge\vartheta^{a}\wedge\vartheta^{b}}_{Nieh-Yan\,\,Term}.\end{split}$ (2.12)
Therefore, the addition of a topological term to the Yang-Mills action naturally produces further contributions that can be identified with well- known topological invariants of gravity [Corichi:2013zza, Nieh:1981ww, Rezende:2009sv, Montesinos:2001ww] [^4] For this reason, with an abuse of notation, we name the topological terms in the same way as their counterparts which will be obtained in the contraction limit.. The question is whether the formal analogy of the (A)dS Yang-Mills model with first-order Einstein-Palatini gravity can be made physical via a contraction of the structure group to Poincaré, and which conditions should be satisfied for that to happen. To this aim, the choice of basis in the (A)dS algebra and its dependence on $\alpha$ play an essential role. In the limit $\alpha\rightarrow 0$, $\mathfrak{g}_{\alpha}$ reduces to the Poincaré algebra but the action does not change in form and it is still well defined, being independent of the parameter, despite the fact that for $\alpha\to 0$, the connection (2.2) and the curvature (2.4) are ill defined. Therefore the action (2.10) will remain gauge invariant only with respect to the Lorentz group, suggesting a consistent identification with some formulation of gravity if the limit is carefully performed.
@@
    the differential identities implied by the vanishing of the exterior covariant derivative of the curvature two-form, $\mathcal{D}_{\Omega}\mathcal{F}=0$ are consistent with the Bianchi identities of the would-be Riemann curvature and torsion;
@@
    the behavior of the gauge fields under the action of the local Lorentz group, in the limit $\alpha\rightarrow 0$, is consistent with their interpretation as tetrad one-form and Lorentz connection;